xydx(2x^2+3y^2-20)dy=0


1
Expert's answer
2021-11-01T19:59:48-0400

"xydx+(2x^2+3y^2-20)dy=0"


integrating factor:

"\mu (y)=y^3"

"xy^4dx+(2x^2+3y^2-20)y^3dy=0"


"f(x,y)=\int y^4xdx=y^4x^2/2+g(y)"


"(y^4x^2\/2+g(y))_y=2y^3x^2+dg(y)\/dy"


"2y^3x^2+dg(y)\/dy=(2x^2+3y^2-20)y^3"


"dg(y)\/dy=(3y^2-20)y^3"


"g(y)=\int(3y^2-20)y^3dy=y^6/2-5y^4"


"f(x,y)=y^6\/2-5y^4+y^4x^2\/2"


"f(x,y)=c_1"


"y^6\/2-5y^4+y^4x^2\/2=c_1"